OpenStudy (anonymous):

Given the equation ,6x-y=98 find x if y = –2.

OpenStudy (gabylovesyou):

substitute the y for -2... 6x - (-2) = 98 turns into 6x + 2 = 98 subtract 2 on both sides 6x + 2 - 2 = 98 - 2 6x = 96 divide both sides by 6 6x/6 = 96 / 6 x = 16 :)